OpenStudy (anonymous):

7. 4(y – 4) = 8 (1 point) –2 2 4 6

OpenStudy (compassionate):

Hello, Distribute 4. 4y - 16 = 8 Add 16 4y = 24 y = 6 IF you found this helpful @Ihateschoolsomuchoml , let me know! :)

OpenStudy (anonymous):

yes i did thank u(:

OpenStudy (solomonzelman):

another approach Divide both sides by 4, this way the 4 on the right is canceled out. 4(y – 4) = 8 y - 4 = 2 +4 +4 y = 6